| <p>18:19 Start or run 7: In beam has been optimzed up to now by Zoran and Claude. ~7E7 particles at injection. From now on, only the injection beamline is optimized by Operators.</p>
<p>Detectors IN (left arm pot =18.86 mm, right arm pot = 19.80 mm)</p>
<p>No Target</p>
<p>Beam is in the ring for ~6 seconds and total cycle time between injections is ~8 seconds</p>
<p>Injection beamline is still in process of optimization by operators</p>
<p>21:30</p>
<p>Merger OK, Tape server OK</p>
<p>ADC data item statistics looks as before. OK</p>
<p>FEE temperatures look OK</p>
<p>System checks OK</p>
<p>Leakage current stable</p>
<p>23:30</p>
<p>Merger OK, Tape server OK</p>
<p>ADC data item statistics looks as before. OK</p>
<p>FEE temperatures look OK</p>
<p>System checks OK</p>
<p>Leakage current stable</p>
<p>01:30</p>
<p>Merger OK, Tape server OK</p>
<p>ADC data item statistics looks as before. OK</p>
<p>FEE temperatures look OK</p>
<p>System checks OK</p>
<p>Leakage current stable</p>
<p>Beam and Electron Coller failure (~00:35); Beam back (~01:08); Cooler back (~01:20)</p>
<p>03:00</p>
<p>Merger OK, Tape server OK</p>
<p>ADC data item statistics looks as before. OK</p>
<p>FEE temperatures look OK</p>
<p>System checks OK</p>
<p>Leakage current stable</p>
<p>Beam and Electron Coller failure (~01:35); Beam back (~01:55); Cooler back (~01:55)</p>
<p>04:40</p>
<p>Merger OK, Tape server OK</p>
<p>ADC data item statistics looks as before. OK</p>
<p>FEE temperatures look OK</p>
<p>System checks OK</p>
<p>Leakage current stable</p>
<p>06:00</p>
<p>Merger OK, Tape server OK</p>
<p>ADC data item statistics looks as before. OK</p>
<p>FEE temperatures look OK</p>
<p>System checks OK</p>
<p>Leakage current stable</p>
<p>07:00</p>
<p>Merger OK, Tape server OK</p>
<p>ADC data item statistics looks as before. OK</p>
<p>FEE temperatures look OK</p>
<p>System checks OK</p>
<p>Leakage current stable</p>
<p> </p>
<p> </p> |